# On the Verification of Control Flow Attestation Evidence

### v2

@@ Abstract.
In remote run-time attestation, including Control Flow Attestation (CFA) and Data Flow Attestation (DFA), a remote Verifier ($\sf{\mathcal{V}rf}$) requests a potentially compromised Prover device ($\sf{\mathcal{P}rv}$) to generate evidence of its execution control flow path (in CFA) and optionally execution data inputs (in DFA). Recent advances in this space also guarantee that $\sf{\mathcal{V}rf}$ eventually receives run-time evidence from $\sf{\mathcal{P}rv}$, even when $\sf{\mathcal{P}rv}$ is fully compromised. Reliable delivery, in theory, enables secure run-time auditing in addition to attestation, in which $\sf{\mathcal{V}rf}$ examines the evidence to identify exploited vulnerabilities and pinpoint/remediate their root cause(s). However, prior work has left $\sf{\mathcal{V}rf}$’s perspective in this security service unexplored, primarily focusing on methods to securely generate authentic run-time evidence on the (potentially compromised) $\sf{\mathcal{P}rv}$. In this work, we argue that run-time attestation/auditing is only effective if $\sf{\mathcal{V}rf}$ can analyze the received evidence. From this premise, we characterize the different types of evidence produced by run-time attestation/auditing architectures by $\sf{\mathcal{V}rf}$’s ability to use them for vulnerability detection/remediation. As a case study, we propose SABRE– a Security Analysis and Binary Repair Engine – that enables $\sf{\mathcal{V}rf}$ to use run-time evidence to detect control flow attacks, to pinpoint specific instructions that corrupted control data, and to automatically generate binary patches to buffer overflow and use-after-free vulnerabilities without source code knowledge.
@@ 1. Introduction
Embedded devices exist at the edge of larger systems, often serving as remotely operated wireless sensors and on-demand actuators. Due to low energy and cost requirements, they are commonly implemented using low-end micro-controller units (MCUs) that lack advanced architectural security features (e.g., virtual memory or memory management units). Consequently, MCUs are more vulnerable and frequently targeted by cyber attacks (Nafees et al., 2023; Kayan et al., 2022).
@@
To address this limitation, RA can be extended into Control Flow Attestation (CFA). In CFA, $\sf{\mathcal{V}rf}$ obtains cryptographic proof of both the installed software image and the exact control flow path followed in the most recent execution of the software (Ammar et al., 2024). This evidence and the respective proof are generated/authenticated by a root of trust (RoT) within $\sf{\mathcal{P}rv}$ that securely records control flow transfer destinations. Therefore, CFA evidence can, in principle, allow $\sf{\mathcal{V}rf}$ to detect both code modification and control-flow hijacking attacks. Recent architectures augment CFA to obtain run-time auditing (Caulfield et al., 2023, 2024a) by guaranteeing the delivery of CFA evidence to $\sf{\mathcal{V}rf}$ even when $\sf{\mathcal{P}rv}$’s software is compromised. See Sec. 2 for details on CFA and run-time auditing RoTs.
@@
We observe that prior work in this area (Ammar et al., 2024) has focused on secure design/implementation of $\sf{\mathcal{P}rv}$ RoTs, often overlooking $\sf{\mathcal{V}rf}$’s role. To our knowledge, no systematic analysis has explored $\sf{\mathcal{V}rf}$’s ability to interpret CFA evidence beyond basic path validation (i.e., determining whether the reported control-flow path is valid). However, in practice, after deeming a path invalid consequently detecting a compromise on $\sf{\mathcal{P}rv}$, $\sf{\mathcal{V}rf}$ would need to use the CFA evidence to (1) identify the root cause of this compromise and (2) implement patches to mitigate it, neither of which has been thoroughly investigated in the literature. These challenges become even more difficulty when $\sf{\mathcal{V}rf}$ does not fully own $\sf{\mathcal{P}rv}$’s software and its executable contains proprietary dependencies/third-party libraries with unavailable source code: a common occurrence in embedded system software chains. Given the prevalence of memory-unsafe languages in embedded software and their various associated exploits ((2023), CISA), we argue that automated methods to identify unknown memory safety vulnerabilities from attested run-time evidence can result in timely detection of otherwise oblivious attacks and elimination of their causes.
@@
While techniques have been proposed for root cause vulnerability analysis in high-end devices (e.g., based on Intel PT traces (Intel, 2015)), none of them can be used to analyze execution traces produced by CFA. In particular, these techniques (Yagemann et al., 2021b, a; Zeng et al., 2022; Liu et al., 2022) require the RoT to provide initial memory snapshots for a remote server to perform a root cause analysis. This requirement is unlikely to be satisfied in the CFA settings due to two reasons. First, this would require $\sf{\mathcal{P}rv}$ to transmit potentially large memory snapshots during CFA, conflicting with the resource-constrained nature of MCUs (Ammar et al., 2024). Second, this snapshot must be stored by $\sf{\mathcal{P}rv}$ during the attested execution. This, in turn, unnecessarily reduces memory available to regular tasks on $\sf{\mathcal{P}rv}$. We expand on these points in Sec. 6. Due to these conflicts, existing root cause analysis frameworks are not directly compatible with CFA.
Our work bridges these gaps with the following contributions:
@@
    We develop a fully functional open-source prototype of SABRE, supporting TI MSP430 and ARM Cortex-M MCU binaries (Caulfield et al., 2024b) and run-time evidence generated by existing open-source run-time attestation RoTs (based on Trusted Execution Environments (TEEs) (Caulfield et al., 2024a) and custom hardware (Caulfield et al., 2023)).
@@ 2.3. Run-time Attestation and Auditing
Run-time Attestation and Auditing are a set of security services that aim to generate and guarantee the delivery of accurate and authentic evidence of a $\sf{\mathcal{P}rv}$’s run-time behavior to a remote $\sf{\mathcal{V}rf}$, thus reflecting to $\sf{\mathcal{V}rf}$ whether any control flow attack has occurred. One type of run-time attestation is Control Flow Attestation (CFA(Nunes et al., 2021b; Sun et al., 2020; Dessouky et al., 2018; Zeitouni et al., 2017; Toffalini et al., 2019; Zhang et al., 2021; Abera et al., 2016; Geden and Rasmussen, 2019; Huo et al., 2020; Neto and Nunes, 2023; Yadav and Ganapathy, 2023; Armanuzzaman et al., 2025), which extends RA to detect run-time attacks in addition to illegal binary modifications. Upon receiving a request to execute some software function along with a RA challenge from $\sf{\mathcal{V}rf}$, $\sf{\mathcal{P}rv}$ executes the requested function, and a RoT in $\sf{\mathcal{P}rv}$ records a trace of this execution into protected memory. In the most fine-grained case, the trace is a control flow log ($\mathcal{C}F_{Log}$) of all control flow transfers that have occurred during the requested execution. After producing $\mathcal{C}F_{Log}$, $\sf{\mathcal{P}rv}$ performs the authenticated RA integrity measurement over the $\mathcal{C}F_{Log}$, the received challenge, and its own program memory to produce a timely and authenticated report. This report attests to the static state of $\sf{\mathcal{P}rv}$’s program binary and the run-time state within this timely execution. When applicable, data inputs – in Data Flow Attestation (DFA(Sun et al., 2020; Nunes et al., 2021a; Dessouky et al., 2018) – or produced outputs (Nunes et al., 2020; Caulfield et al., 2022) can also be included in this proof to bind obtained results to the proper execution of respective software functions on expected inputs.
Existing CFA techniques use either (1) binary instrumentation along with TEE support; or (2) custom hardware modifications to generate $\mathcal{C}F_{Log}$ by detecting and saving each branch destination to a dedicated and protected memory region. For techniques that use binary instrumentation, a pre-processing phase modifies the binary so that all branch instructions (e.g., jumps, returns, calls, etc.) are prepended with additional calls to a TEE-protected trusted code. Once called, the trusted code appends $\mathcal{C}F_{Log}$ to update the trace with the current branch destination. In hardware-based techniques, custom hardware interfaces with the MCU core to detect branches and records their destinations to a reserved memory region. Recent work (Caulfield et al., 2023, 2024a) has also created mechanisms to implement reliable communication with $\sf{\mathcal{V}rf}$ as a part of the CFA RoT, enabling run-time auditing by ensuring evidence delivery even in the presence of malware that infects $\sf{\mathcal{P}rv}$ and refuses to deliver evidence about its actions/presence.
@@ 3.1. Evidence Types
The various realizations of CFA RoTs (see Sec. 2) create and manage $\mathcal{C}F_{Log}$ differently, presenting trade-offs between the storage/communication cost and ease of verification/analysis by $\sf{\mathcal{V}rf}$. We classify existing methods in three types: hash-based evidence (E1), verbatim evidence (E2), and hybrid evidence (E3).
@@
Hash-based Evidence (E1). As control flow transfers occur, branch destination addresses are accumulated into a single hash chain. This is performed by first accumulating the current destination address with the previous hash and then hashing the accumulated result (Abera et al., 2016; Dessouky et al., 2017; Huo et al., 2020; Zeitouni et al., 2017), as exemplified in Fig. 2.a. After the first branch occurs, its destination 0xe004 is accumulated into the hash. Since this is the first branch destination, it is accumulated with an initial constant to obtain $H_{0}$. Upon the second branch instruction, its destination address 0xf000 is accumulated with $H_{0}$ to obtain $H_{1}$. This process continues until the end of the requested execution generating a unique value that represents the path. Although simple control flows produce a single hash in E1, programs with more complex control flows (including those with nested loops and branching) might require a set of hash digests (Abera et al., 2016; Zeitouni et al., 2017). This approach yields minimal storage/transmission overheads (i.e., the fixed size of a hash digest). However, $\sf{\mathcal{V}rf}$ must perform its analysis solely based on the final hash digest and the executable binary.
Verbatim Evidence (E2). On the other side of the spectrum, more recent CFA techniques advocate for storing and transmitting $\mathcal{C}F_{Log}$ in its entirety (Nunes et al., 2021b; Dessouky et al., 2018; Toffalini et al., 2019; Caulfield et al., 2023; Zhang et al., 2021; Caulfield et al., 2024a). As depicted in Fig. 2.b, the full sequence of destination addresses is maintained in a buffer. Therefore, this approach requires further (lossless) compression or is limited to small and self-contained operations to avoid filling the dedicated storage too quickly. A common strategy is to log simple loops (those without internal branches) with their address once and a count/condition pertaining to the number of iterations performed (Caulfield et al., 2023; Nunes et al., 2021b; Zhang et al., 2021; Caulfield et al., 2024a). Another reduction is obtained by recording statically defined branch destinations as a single bit (Dessouky et al., 2018). Since static destinations can be determined from the binary alone, branch instructions with statically defined destinations can be recorded as a single bit (1/0) to determine if the branch was taken or not. Branches with dynamically defined destinations (e.g., returns and indirect calls) are still recorded with their full addresses. If $\mathcal{C}F_{Log}$ remains larger than available storage despite these optimizations, it can be split into a set of multiple fixed-size slices and streamed to $\sf{\mathcal{V}rf}$, as suggested in (Toffalini et al., 2019; Caulfield et al., 2023, 2024a). Despite the aforementioned challenges, E2 provides $\sf{\mathcal{V}rf}$ with full control flow path information for analysis.
@@ 3.2. Detection & Remediation via E1, E2 & E3
Consider an attack detection strategy in which $\sf{\mathcal{V}rf}$ constructs the CFG of the executable binary, i.e., as suggested in (Abera et al., 2016; Caulfield et al., 2023; Dessouky et al., 2017; Zeitouni et al., 2017; Toffalini et al., 2019; Huo et al., 2020; Zhang et al., 2021). An initial set of invalid paths is defined as forward edges that do not exist in the statically constructed CFG. As valid indirect targets (e.g., return addresses or indirect calls) are over-approximated using the static CFG (Burow et al., 2019), $\sf{\mathcal{V}rf}$ may also use $\mathcal{C}F_{Log}$ along with the program’s binary to emulate the execution of the program locally. In this process, a Shadow Stack (SS) can be emulated locally to verify backward edges (return destinations) and data flow (DF) analysis can be employed to determine a set of valid call sites for each indirect call, updating the CFG accordingly. We observe feasibility of this verification strategy varies greatly depending on evidence types E1, E2, and E3.
Figure 3. SABRE submodules and internal workflow
@@ 4. SABRE
SABRE showcases $\mathcal{C}F_{Log}$-based root cause analysis and remediation with verbatim control flow traces (E2) and without the source-code of the attested program. Fig. 3 shows SABRE workflow consisting of five internal modules: Static Analysis, Path Verifier, Exploit Locator, Patch Generator, and Patch Validator. SABRE takes as input an application binary (App.elf) and a verbatim $\mathcal{C}F_{Log}$ (E2 evidence type). The notation used in the rest of the paper is summarized in Table 1.

### v3

@@ Abstract.
In run-time attestation schemes, including Control Flow Attestation (CFA) and Data Flow Attestation (DFA), a remote Verifier ($\sf{\mathcal{V}rf}$) requests a potentially compromised Prover device ($\sf{\mathcal{P}rv}$) to generate evidence of its execution control flow path (in CFA) and optionally execution data inputs (in DFA). Recent advances in this space also guarantee that $\sf{\mathcal{V}rf}$ eventually receives run-time evidence from $\sf{\mathcal{P}rv}$, even when $\sf{\mathcal{P}rv}$ is fully compromised. Reliable delivery, in theory, enables run-time auditing in addition to attestation, allowing $\sf{\mathcal{V}rf}$ to examine run-time compromise traces to pinpoint/remediate attack root causes. However, $\sf{\mathcal{V}rf}$’s perspective in this security service remains unexplored, with most prior work focusing on the secure generation of authentic run-time evidence on $\sf{\mathcal{P}rv}$.
[…]
journalyear: 2025copyright: acmlicensedconference: 18th ACM Conference on Security and Privacy in Wireless and Mobile Networks; June 30-July 3, 2025; Arlington, VA, USAbooktitle: 18th ACM Conference on Security and Privacy in Wireless and Mobile Networks (WiSec 2025), June 30-July 3, 2025, Arlington, VA, USAdoi: 10.1145/3734477.3734710isbn: 979-8-4007-1530-3/2025/06
@@ 1. Introduction
Embedded devices exist at the edge of larger systems, often serving as remotely operated sensors and on-demand actuators. Due to low energy and cost requirements, they are commonly implemented using low-end micro-controller units (MCUs) that lack advanced architectural security features (e.g., virtual memory or memory management units). Consequently, MCUs are more vulnerable and frequently targeted by cyber attacks (Nafees et al., 2023; Kayan et al., 2022).
@@
To address this limitation, RA can be extended into Control Flow Attestation (CFA). In CFA, $\sf{\mathcal{V}rf}$ obtains cryptographic proof of both the installed software image and the exact control flow path followed in the most recent execution of the software (Ammar et al., 2024). This evidence and the respective proof are generated/authenticated by a root of trust (RoT) within $\sf{\mathcal{P}rv}$ that securely records control flow transfer destinations. Therefore, CFA evidence can, in principle, allow $\sf{\mathcal{V}rf}$ to detect both code modification and control-flow hijacking attacks. Recent architectures augment CFA to obtain run-time auditing (Caulfield and otehrs, 2023; Caulfield et al., 2024b) by guaranteeing the delivery of CFA evidence to $\sf{\mathcal{V}rf}$ even when $\sf{\mathcal{P}rv}$’s software is compromised. See Sec. 2 for details on CFA and run-time auditing RoTs.
@@
We observe that prior work in this area (Ammar et al., 2024) has focused on secure design/implementation of $\sf{\mathcal{P}rv}$ RoTs, often overlooking $\sf{\mathcal{V}rf}$’s role. To our knowledge, no systematic analysis has explored $\sf{\mathcal{V}rf}$’s ability to interpret CFA evidence beyond basic path validation (i.e., determining whether the reported control-flow path is valid). However, in practice, after deeming a path invalid (thus $\sf{\mathcal{P}rv}$ compromise), $\sf{\mathcal{V}rf}$ would need to use CFA evidence to (1) identify the root cause of this compromise and (2) implement patches to mitigate it. Neither of these has been thoroughly investigated in the literature. The challenge increases when $\sf{\mathcal{P}rv}$’s software contains proprietary dependencies/third-party libraries with unavailable source code: a common occurrence in embedded system software chains.
@@
Given the prevalence of memory-unsafe languages in embedded software and their various associated exploits ((2023), CISA), we argue that automated methods to identify unknown memory safety vulnerabilities from attested run-time evidence can result in timely detection of otherwise oblivious attacks and elimination of their root causes. While techniques have been proposed for root cause vulnerability analysis in high-end devices (e.g., based on Intel PT traces (Intel, 2015)), none of them can be used to analyze execution traces produced by CFA. In particular, these techniques (Yagemann et al., 2021a, b; Zeng et al., 2022; Liu et al., 2022) require initial memory snapshots for a remote server to perform a root cause analysis. This requirement is unlikely to be satisfied in the CFA settings for two reasons. First, this would require $\sf{\mathcal{P}rv}$ to transmit potentially large memory snapshots during CFA, conflicting with the resource-constrained nature of MCUs (Ammar et al., 2024). Second, this snapshot must be stored by $\sf{\mathcal{P}rv}$ during the attested execution. This, in turn, unnecessarily reduces memory available to regular tasks on $\sf{\mathcal{P}rv}$. We expand on these points in Sec. 6. Due to these conflicts, existing root cause analysis frameworks are not directly compatible with CFA. Our work bridges these gaps with the following contributions:
@@
    We develop an open-source prototype of SABRE, supporting TI MSP430 and ARM Cortex-M MCU binaries (Caulfield et al., 2025) and run-time evidence generated by existing open-source run-time attestation RoTs (based on Trusted Execution Environments (TEEs) (Caulfield et al., 2024b) and custom hardware (Caulfield and otehrs, 2023)).
@@ 2.3. Run-time Attestation and Auditing
Run-time Attestation and Auditing are a set of security services that aim to generate and guarantee the delivery of accurate and authentic evidence of a $\sf{\mathcal{P}rv}$’s run-time behavior to a remote $\sf{\mathcal{V}rf}$, thus reflecting to $\sf{\mathcal{V}rf}$ whether any control flow attack has occurred. One type of run-time attestation is Control Flow Attestation (CFA(Nunes et al., 2021b; Sun et al., 2020; Dessouky et al., 2018; Zeitouni et al., 2017; Toffalini et al., 2019; Zhang et al., 2021; Abera et al., 2016; Geden and Rasmussen, 2019; Neto and Nunes, 2023; Yadav and Ganapathy, 2023; Armanuzzaman et al., 2025), which extends RA to detect run-time attacks in addition to illegal binary modifications. Upon receiving a request to execute some software function along with a RA challenge from $\sf{\mathcal{V}rf}$, $\sf{\mathcal{P}rv}$ executes the requested function, and a RoT in $\sf{\mathcal{P}rv}$ records a trace of this execution into protected memory. In the most fine-grained case, the trace is a control flow log ($\mathcal{C}F_{Log}$) of all control flow transfers that have occurred during the requested execution. After producing $\mathcal{C}F_{Log}$, $\sf{\mathcal{P}rv}$ performs the authenticated RA integrity measurement over the $\mathcal{C}F_{Log}$, the received challenge, and its own program memory to produce a timely and authenticated report. This report attests to the static state of $\sf{\mathcal{P}rv}$’s program binary and the run-time state within this timely execution. When applicable, data inputs – in Data Flow Attestation (DFA(Sun et al., 2020; Nunes et al., 2021a; Dessouky et al., 2018) – or produced outputs (Nunes et al., 2020; Caulfield et al., 2022) can also be included in this proof to bind obtained results to the proper execution of respective software functions on expected inputs.
Existing CFA techniques use either (1) binary instrumentation along with TEE support; or (2) custom hardware modifications to generate $\mathcal{C}F_{Log}$ by detecting and saving each branch destination to a dedicated and protected memory region. For techniques that use binary instrumentation, a pre-processing phase modifies the binary so that all branch instructions (e.g., jumps, returns, calls, etc.) are prepended with additional calls to a TEE-protected trusted code. Once called, the trusted code appends $\mathcal{C}F_{Log}$ to update the trace with the current branch destination. In hardware-based techniques, custom hardware interfaces with the MCU core to detect branches and records their destinations to a reserved memory region. Recent work (Caulfield and otehrs, 2023; Caulfield et al., 2024b) has also created mechanisms to implement reliable communication with $\sf{\mathcal{V}rf}$ as a part of the CFA RoT, enabling run-time auditing by ensuring evidence delivery even in the presence of malware that infects $\sf{\mathcal{P}rv}$ and refuses to deliver evidence about its actions/presence.
@@ 3.1. Evidence Types
The various realizations of CFA RoTs (see Sec. 2) create and manage $\mathcal{C}F_{Log}$ differently, presenting trade-offs between the storage/communication cost and ease of verification/analysis by $\sf{\mathcal{V}rf}$. We classify existing methods in three types: hash-based evidence (E1), verbatim evidence (E2), and hybrid evidence (E3). Compared to Ammar et al. (Ammar et al., 2024), we further offer a more detailed analysis of each type’s suitability for advanced analysis.
@@
Hash-based Evidence (E1). As control flow transfers occur, branch destination addresses are accumulated into a single hash chain. This is performed by first accumulating the current destination address with the previous hash and then hashing the accumulated result (Abera et al., 2016; Dessouky et al., 2017; Zeitouni et al., 2017), as exemplified in Fig. 2.a. After the first branch occurs, its destination 0xe004 is accumulated into the hash. Since this is the first branch destination, it is accumulated with an initial constant to obtain $H_{0}$. Upon the second branch instruction, its destination address 0xf000 is accumulated with $H_{0}$ to obtain $H_{1}$. This process continues until the end of the requested execution generating a unique value that represents the path. Although simple control flows produce a single hash in E1, programs with more complex control flows (including those with nested loops and branching) might require a set of hash digests (Abera et al., 2016; Zeitouni et al., 2017). This approach yields minimal storage/transmission overheads (i.e., the fixed size of a hash digest). However, $\sf{\mathcal{V}rf}$ must perform its analysis solely based on the final hash digest and the executable binary.
Verbatim Evidence (E2). On the other side of the spectrum, more recent CFA techniques advocate for storing and transmitting $\mathcal{C}F_{Log}$ in its entirety (Nunes et al., 2021b; Dessouky et al., 2018; Toffalini et al., 2019; Caulfield and otehrs, 2023; Zhang et al., 2021; Caulfield et al., 2024b). As depicted in Fig. 2.b, the full sequence of destination addresses is maintained in a buffer. Therefore, this approach requires further (lossless) compression or is limited to small and self-contained operations to avoid filling the dedicated storage too quickly. A common strategy is to log simple loops (those without internal branches) with their address once and a count/condition pertaining to the number of iterations performed (Caulfield and otehrs, 2023; Nunes et al., 2021b; Zhang et al., 2021; Caulfield et al., 2024b). Another reduction is obtained by recording statically defined branch destinations as a single bit (Dessouky et al., 2018). Since static destinations can be determined from the binary alone, branch instructions with statically defined destinations can be recorded as a single bit (1/0) to determine if the branch was taken or not. Branches with dynamically defined destinations (e.g., returns and indirect calls) are still recorded with their full addresses. If $\mathcal{C}F_{Log}$ remains larger than available storage despite these optimizations, it can be split into a set of multiple fixed-size slices and streamed to $\sf{\mathcal{V}rf}$, as suggested in (Toffalini et al., 2019; Caulfield and otehrs, 2023; Caulfield et al., 2024b). Despite the aforementioned challenges, E2 provides $\sf{\mathcal{V}rf}$ with full control flow path information for analysis.
@@ 3.2. Detection & Remediation via E1, E2 & E3
Figure 3. SABRE sub-modules and internal workflow
Consider an attack detection strategy in which $\sf{\mathcal{V}rf}$ constructs the CFG of the executable binary, i.e., as suggested in (Abera et al., 2016; Caulfield and otehrs, 2023; Dessouky et al., 2017; Zeitouni et al., 2017; Toffalini et al., 2019; Zhang et al., 2021). An initial set of invalid paths is defined as forward edges that do not exist in the statically constructed CFG. As valid indirect targets (e.g., return addresses or indirect calls) are over-approximated using the static CFG (Burow et al., 2019), $\sf{\mathcal{V}rf}$ may also use $\mathcal{C}F_{Log}$ along with the program’s binary to emulate the execution of the program locally. In this process, a Shadow Stack (SS) can be emulated locally to verify backward edges (return destinations) and data flow (DF) analysis can be employed to determine a set of valid call sites for each indirect call, updating the CFG accordingly. We observe feasibility of this verification strategy varies greatly depending on evidence types E1, E2, and E3.
@@ 4. SABRE
SABRE showcases $\mathcal{C}F_{Log}$-based root cause analysis and remediation with verbatim control flow traces (E2) and without the source-code of the attested program. Fig. 3 shows SABRE workflow consisting of five internal modules: Static Analysis, Path Verifier, Exploit Locator, Patch Generator, and Patch Validator. SABRE takes as input an application binary (App.elf) and a verbatim $\mathcal{C}F_{Log}$ (E2 evidence type). The notation used in this work is summarized in Table 1.
4.1. Static Analysis and Path Verifier
Figure 4. CFG and $\mathcal{C}F_{Log}$
Figure 5. Backward Traversal for a ret
Figure 6. Backward Traversal for an call